YoursTruly:

Solve the system by substitution. Show your work. y=-x^2 y=2x+4

AshtynThatDrummerGirl:

y = -0^2 y = 0 0 = 2x + 4 -4 -4 -4 = 2x /2 /2 -2 = x (-2, 0) is your answer

YoursTruly:

tyty

AshtynThatDrummerGirl:

ywyw

YoursTruly:

x'D